OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -9 or x + 5 12 A) A number line with an open circle on -6, shading to the left, and a closed circle on 7, shading to the right. B) A number line with an open circle on -6, a closed circle on 7, and shading in between. C) A number line with a closed circle on -6, shading to the left, and an open circle on 7, shading to the right. d) A number line with a closed circle on -6, an open circle on 7, and shading in between.

OpenStudy (bahrom7893):

x - 3 < -9, or x - 3 + 3 < -9+3 x < -6 What's the other one? you forgot the inequality sign.

OpenStudy (anonymous):

5 is greatethan or equal to 12.

OpenStudy (bahrom7893):

x + 5>=12 x + 5 - 5 >= 12-5 x >= 7

OpenStudy (bahrom7893):